OpenStudy (anonymous):

The graph of y = x2 is shown below. Using complete sentences, explain how the graph of y = 2(x + 3)2 – 4 would differ from this parabola?

OpenStudy (anonymous):

its the graph of y=x^2 shifted left 3 units, stretched horizontally by a factor of 2, and shifted down 4 units

OpenStudy (anonymous):

stretched vertically by factor of 2****

OpenStudy (anonymous):



OpenStudy (anonymous):

y=x^2 graph has axis of symmetry x = 0 ; y = 2(x + 3)2 – 4 graph has axis of symmetry x = -3

OpenStudy (anonymous):

mapple -> graph